onsider the single stage amplifier, values of h-parameters are hie=X1.X2 kΩ, hfe=(150-X1), hre=X1. X2˟10-4 and hoe=50µA/V. Rs=X1X2 Ω, Calculate Av, Ai, Ri, Zo


1
Expert's answer
2021-05-10T02:42:17-0400

"Av=\dfrac{h_{f_e}}{h_{re}h_{fe}-h_{ve}(R_s+h_{ie})}

\\[9pt]

 =\dfrac{149}{1.3\times 10^{-4}\times 149-50\times 10^{-6}\times (0.013+1.3)}"


Av=7718.46 


"Ai=-h_{fe}=-149"


As we know ,


"R_{in}=h_{ie}-\dfrac{h_{fe}h_{re}}{h_{oe}}
\\[9pt]


 =1.3-\dfrac{149\times 1.3\times 10^{-4}}{50\times 10^{-6}}"


"R_{in}=0.912 K\Omega"


"Z_o=\dfrac{1}{h_{oe}-\frac{h_{fe}h_{re}}{R_s+h_{ie}}}"


   "=\dfrac{1}{50\times 10^{-6}-\frac{1.49\times 1.3\times 10^{-4}}{0.013+1.3\times 10^3}}
\\[9pt]
 =28.489K\Omega"